# Weak Memory Model Formalisms: Introduction and Survey

### v1

@@ 3.3. Software Memory Models
Sun Microsystems released the Java programming language in 1996, with multi-threading and shared-memory behaviour informally described in [137].
This was immediately followed by multiple independent efforts to formalise the memory model of Java [43; 132; 148]. There is a formulation that explicitly used local working memories [58], and a more abstract formulation with event structures [76]. In 1999, however, the original prose memory model turned out to contain flaws [238; 239]. Subsequent work [203; 206; 252; 296; 297; 44; 207] attempted to revise the Java memory model, culminating in [208], which describes the new Java memory model as part of Java 5.0.
After the formulation in [208] became accepted as the official Java model, a few alternative specifications continued to emerge [298; 154; 195; 53]. On the compiler-related side, Ševčík and Aspinall [258] identified common optimisations performed Java compilers that do not comply with the Java memory model.
@@
Boehm and others worked as part of the International Organisation for Standardisation (ISO) on the formal semantics of concurrent C and C++ programs [55; 56]. The resulting 2011 ISO standards of the C and C++ programming languages were the first time that multi-threading in both languages was formally defined, and the memory model defined by these ISO standards is commonly referred to as the C memory model (or C11 in the literature, owing to its inception in 2011). The definition of the C memory model was then consolidated into a formal axiomatic formulation by Batty et al. [50], and later into a formal operational formulation by Nienhuis et al. [221].
The prominence of the C language and its memory model has drawn much attention in the context of verification, and many recent publications are devoted to program logics [295], model checkers [173], or other problems [266; 265; 86]; more of these are covered in Section 8.
Software memory models are often subject to more potential issues, due to the wider range of transformations open to a compiler, on top of the weak memory behaviours on the target hardware. These issues include the “out of thin air” (see Example 3.3) and “read from untaken branch” behaviours. These have been particularly addressed in the context of C, and the question about how to forbid out-of-thin air (and similar) behaviours without overly constraining standard compiler optimisations is a subject of ongoing debate in the standardisation committees and the academic community [57; 48; 155; 232; 226; 79]. The most recent C reference explicitly gives a version of out-of-thin-air behaviour, stating that it is a theoretically allowed behaviour, but with a caveat that no compiler should ever implement it.
[ht]
With the memory model of Power lacking multicopy atomicity and hence being harder to handle (see Section 3.2), there have been works that focus on the compilation from C to Power while respecting the memory models of both ends [49; 253; 183], and more general compiler optimisations under the C memory model [233; 109; 126].
@@
Alglave et al. [32] provide a formalisation of the memory model of the Linux kernel. This model is to further codify the complex interaction of the wide variety of hardware memory models and the C language memory model, supporting some of the common patterns specific to Linux kernel programming. The formalisation involved interaction with the developer community, and the effort helped identify and resolve several issues, especially regarding locks and the read-copy-update (RCU) paradigm. As such this work represents a strong argument for the practicality of involving formalisms in non-trivial software development.
@@
Every programming language that supports shared-memory concurrency has interplay with weak memory models. For programming languages such as JavaScript [135; 290], OCaml [114], OpenCL [167] and Rust [102], there have been works that study their own memory models, or how they are affected by the weak memory models of the languages to which they may be compiled. A potential simplification for this effort for many languages is via a memory model for an intermediate language such as LLVM [78; 186], and through the investigation of general properties related to safe optimisation strategies [136].
@@ 4. Introduction to Operational Formalisms
Before giving a general survey of the field we provide an example formalisation: a simplification of the formalisation of x86 given by the influential work by Owens et al. [227; 260]. This formalisation is operational, meaning described by way of individual steps of the system, and as such is often easier to follow than other types of semantics [159]. We also choose it as it is based around a microarchitectural feature, and as such provides a contrast to axiomatic descriptions, which purposefully abstract away from such considerations. In this formalisation, weak memory effects (specifically, loads appearing to take effect before earlier stores) emerge by the use of a write buffer in each core (acting similarly to a pipeline (Section 2.1) or as part of the cache system [212]). We first give an operational semantics for a standard sequential language, as if instructions were executed in the order specified by the program text; after extending that semantics with extra rules covering write buffers the weaker behaviours of x86 (Total Store Order) become apparent. More complex formalisations based on microarchitectural mechanisms are outlined in Section 4.4; see also Section 6.1.
@@ 4.3. Write Buffers for x86
We now augment the system of Sections 4.1 and 4.2 with an explicit write buffer for each core, following [227]. A write buffer, of type , is a list of pairs , where is a shared variable and is a value being written to , representing a pending write of to . The head (left-hand side) of the write buffer is the oldest write, with new writes being appended to the end (right-hand side) of the write buffer. A core now contains a write buffer, in addition to the local state and command; the system-level structure is otherwise unchanged.
@@ 4.4. Operational Descriptions of More Complex Models
Below, we outline how some of these more complex aspects of weak memory may be approached through an explicit model of a hardware pipeline as described in Section 2.1. However, a focus on microarchitectural details may become cumbersome to work with formally. For this reason many researchers focus on formal representations that abstract away from microarchitectural mechanisms, and focus instead on key relationships between operations across multiple processes in the system. This is typically referred to as an axiomatic approach, to which we turn our attention in Section 5. Operationally one of the best-known approaches is the Promising Semantics [165; 187; 241], as discussed in Section 6.2, which combines an operational semantics with global axioms.
@@
In this case we assume all instructions, not just writes, are fetched into the pipeline (generalising the buffer), similarly to (x86-write-buf). Provided that there are no writes to (of the form ), reads of (of the form ), or fences in the section of the pipeline preceding the write in question (), then the write can be issued. A similar rule can be constructed for reads. These constraints encode those of Definition 2.1, modified by fences as in Example 2.2, and allow the weaker behaviours shown in Example 2.1. A more complete operational description of pipelines that conform to the Arm memory model, and includes branches and jumps, is given in [90].
@@
The release/acquire concepts [127], as introduced in (1) and Example 2.2, can also be described operationally. Continuing the generalisation of write buffers to pipelines, an operation with release semantics can be executed only when at the head of the pipeline; i.e., when all other prior operations (w.r.t the program order) have completed. A load with acquire semantics prevents later memory operations from completing earlier; hence we extend the constraints of (Arm-write-mem) to capture this possibility.
Rule (release-write) is similar to (x86-write-mem) in that a release write is issued only when at the head of the pipeline. Rule (RA-write-mem) extends (Arm-write-mem) with an extra condition, which we highlight with a box. This additional constraint means that an earlier acquire load of any variable prevents a later write from proceeding. For the stronger form of release/acquire consistency in [127] an acquire load additionally requires the absence of release stores in the earlier pipeline, which can be described syntactically similarly to the boxed constraint of (RA-write-mem).
@@
Instruction types such as swap, compare-and-swap, fetch-and-increment, etc., combine read and write operations atomically. These can be described using the concept of a global lock which prevents a core from performing a memory operation while another core holds the lock [261]. For instance, the atomic swap instruction, , updates the shared variable to hold the value in register , and updates to the initial value of ; this instruction can be defined as follows.
@@ 5.1. Fundamental Relations
| po | program order | Def. 5.2 | sb | sequenced-before | [50] |
@@
| co | coherence order | Def. 5.5 | mo | modification order | [50] |
| ws | write serialization | [22] | |||
@@
On the name “reads-from”, here we follow the recent work by Arm [27], where a typical pair means “ reads from ”, with occurring before . Some may find it more intuitive to think of the relation rf as writes-to. Indeed, means that writes to the value read by .
The relations po, co and rf are the “base relations”. The set of memory events and these base relations constitute an “execution graph”, also known as a “candidate execution” [27].
@@
The extended coherence order – defined to be the transitive closure of com [183].
@@ 5.2. Axiomatic Definition of Sequential Consistency
A memory model is said to satisfy the coherence property if all possible executions of any execution graph is coherent. The coherence property is also known as the uniproc property [22], SC per location [36], SC per variable [26] or internal visibility [27].
@@ 5.4. Axiomatic Descriptions of More Complex Models
For illustrative purposes, we have presented how axiomatic formalisms can define the straightforward SC memory model and a simplified version of TSO. In practice, axiomatic formalisms have been used to specify a wide range of behaviours related to more complex memory models and instruction types, e.g., [36; 240; 27], as surveyed in Section 6.3. Below we give an indication of how these more complex problems are tackled axiomatically.
@@
Now, a pair belongs to the relation precisely when and are interleaved by a write issued by another thread. In other words, means that and are not atomic. Hence, to ensure that every pair in rmw is atomic, we require that no pair in rmw is in , that is, an execution graph enforces the RMW instructions if . Further details and mechanised tool support can be found in the online repository of the Herd tool [39; 27].
@@ 6.1. Mechanistic Formalisms
Mechanistic formalisms describe the effects of weak memory models with structures that operate in the manner of hardware mechanisms. The best-known of such mechanisms is the local write buffers elucidated in Section 4. Local write buffers describe the behaviour of TSO [263], on which the x86 memory model is based [227; 255; 260]. As local write buffers are relatively straightforward to encode, they often serve as the foundation for verification under TSO. Namely, when each write buffer is modelled as a separate process, one can analyse the entire system under Sequential Consistency using existing tools [68; 280; 281; 292; 282; 190; 191; 192; 273].
The mechanistic formalism of Boudol and Petri [63] is also based on write buffers that are not local but potentially hierarchical, making this formalism capable of expressing behaviours weaker than those of TSO. Another recent variant of write buffers is read buffers, proposed by Abdulla et al. [4] to improve various aspects of verification under TSO. Meanwhile, two other major hardware platforms – Power [254] and Arm [122; 240] – also have their own mechanistic formulations, albeit much more complex than x86-TSO and harder to encode for verification tools. A semantics based directly on pipelined and speculative execution, as outlined in Section 2.1, has been developed for Arm, RISC-V and x86 processors [90].
@@ 6.2. General Operational Formalisms: Timestamped Messages, Promises and Views
Hardware-like mechanisms are insufficient for describing the behaviours of compilers – the source of complexity for the C memory model especially. In order to reason about the C memory model operationally, many formulations make use of the notions of timestamped messages and views. Under weak memory models, a write does not become visible to other threads immediately. Thus, each write can be seen as a message that is broadcast by the issuing process, and it is up to the memory model that dictates when this message becomes visible to which other processes. Each process has a view of a subset of all the messages that have been sent out, and each message has a timestamp attached. By looking at the timestamps, a process updates its view according to the memory model. For instance, instead of just , a write-event is now a pair , where is the time at which the event took place. The timestamp can be some future time, meaning that any subsequent read of with return-value must occur at some later time , i.e., . An early use of timestamps is provided in [234]; we look at some other prominent formalisations in this vein below.
@@
In this framework, introduced in Kang et al. [165], each timestamped message is a promise to fulfil a later write. This semantics was later simplified by Lee et al. [187], and adapted to Arm and RISC-V by Pulte et al. [241]. The semantics has been used as the basis for verification of non-trivial systems [277].
A characteristic of the Promising Semantics is that it is a multi-execution formalism [218], as opposed to a single-execution formalism. In a single-execution formalism (e.g., that of Section 4), one behaviour, e.g., a trace of memory operations, corresponds to one execution of the code. In the Promising Semantics a single behaviour comprises more information that just the sequence memory events, and this extra information also covers multiple possible executions of the code. The multi-execution characteristic can also be applied to partial-order-based representations (see Section 6.4).
@@
Formalisms with timestamped messages and similar notions have since been developed on several fronts, for example, Doherty et al. [111], with subsequent works leading to program logics [96; 295; 110; 98; 99; 45; 291]. Lahav and Boker [178] introduced the notion of potentials. In contrast to timestamped messages which reason about stores that have occurred in the past, potentials reason about possible stores in the future.
Timestamped messages, views, and potentials enable general frameworks for handling a wide range of weak memory effects, including the lack of multicopy atomicity (Section 2.4) and out-of-thin air behaviour (Section 3.3). This flexibility is not without disadvantages, however: these formalisms are perhaps the most complex of those surveyed, with the details of even small examples such as the store buffering litmus test being difficult to reason about [218].

### v2

@@ 3.3. Software Memory Models
Sun Microsystems released the Java programming language in 1996, with multi-threading and shared-memory behaviour informally described in [Gosling et al., 1996].
This was immediately followed by multiple independent efforts to formalise the memory model of Java [Attali et al., 1998; Gontmakher and Schuster, 1998; Higham and Kawash, 1998]. There is a formulation that explicitly used local working memories [Börger and Schulte, 1999], and a more abstract formulation with event structures [Cenciarelli et al., 1999]. In 1999, however, the original prose memory model turned out to contain flaws [Pugh, 1999, 2000]. Subsequent work [Maessen et al., 2000; Manson and Pugh, 2001; Roychoudhury and Mitra, 2002; Yang et al., 2001, 2002; Awhad and Wallace, 2003; Manson and Pugh, 2004] attempted to revise the Java memory model, culminating in [Manson et al., 2005], which describes the new Java memory model as part of Java 5.0.
After the formulation in [Manson et al., 2005] became accepted as the official Java model, a few alternative specifications continued to emerge [Yang et al., 2005; Jagadeesan et al., 2010; Lochbihler, 2013; Bender and Palsberg, 2019]. On the compiler-related side, Ševčík and Aspinall [2008] identified common optimisations performed Java compilers that do not comply with the Java memory model.
@@
Boehm and others worked as part of the International Organisation for Standardisation (ISO) on the formal semantics of concurrent C and C++ programs [Boehm, 2007; Boehm and Adve, 2008]. The resulting 2011 ISO standards of the C and C++ programming languages were the first time that multi-threading in both languages was formally defined, and the memory model defined by these ISO standards is commonly referred to as the C memory model (or C11 in the literature, owing to its inception in 2011). The definition of the C memory model was then consolidated into a formal axiomatic formulation by Batty et al. [2011], and later into a formal operational formulation by Nienhuis et al. [2016].
The prominence of the C language and its memory model has drawn much attention in the context of verification, and many recent publications are devoted to program logics [Wright et al., 2021], model checkers [Kokologiannakis et al., 2018], or other problems [Singh et al., 2021, 2022; Colvin, 2022]; more of these are covered in Section 8.
Software memory models are often subject to more potential issues, due to the wider range of transformations open to a compiler, on top of the weak memory behaviours on the target hardware. These issues include the “out of thin air” (see Example 3.3) and “read from untaken branch” behaviours. These have been particularly addressed in the context of C, and the question about how to forbid out-of-thin air (and similar) behaviours without overly constraining standard compiler optimisations is a subject of ongoing debate in the standardisation committees and the academic community [Boehm and Demsky, 2014; Batty et al., 2015; Jeffrey and Riely, 2016; Pichon-Pharabod and Sewell, 2016; Ou and Demsky, 2018; Chakraborty and Vafeiadis, 2019]. The most recent C reference explicitly gives a version of out-of-thin-air behaviour, stating that it is a theoretically allowed behaviour, but with a caveat that no compiler should ever implement it.
With the memory model of Power lacking multicopy atomicity and hence being harder to handle (see Section 3.2), there have been works that focus on the compilation from C to Power while respecting the memory models of both ends [Batty et al., 2012; Sarkar et al., 2012; Lahav et al., 2017], and more general compiler optimisations under the C memory model [Podkopaev et al., 2019; Dodds et al., 2018; Geeson and Smith, 2024].
@@
Alglave et al. [Alglave et al., 2018] provide a formalisation of the memory model of the Linux kernel. This model is to further codify the complex interaction of the wide variety of hardware memory models and the C language memory model, supporting some of the common patterns specific to Linux kernel programming. The formalisation involved interaction with the developer community, and the effort helped identify and resolve several issues, especially regarding locks and the read-copy-update (RCU) paradigm. As such this work represents a strong argument for the practicality of involving formalisms in non-trivial software development.
@@
Every programming language that supports shared-memory concurrency has interplay with weak memory models. For programming languages such as JavaScript [Gopalakrishnan and Verbrugge, 2020; Watt et al., 2020], OCaml [Dolan et al., 2018], OpenCL [Khronos OpenCL Working Group, 2020] and Rust [Dang et al., 2019], there have been works that study their own memory models, or how they are affected by the weak memory models of the languages to which they may be compiled. A potential simplification for this effort for many languages is via a memory model for an intermediate language such as LLVM [Chakraborty and Vafeiadis, 2017; Lee et al., 2023], and through the investigation of general properties related to safe optimisation strategies [Gopalakrishnan et al., 2023].
@@ 4. Introduction to Operational Formalisms
Before giving a general survey of the field we provide an example formalisation: a simplification of the formalisation of x86 given by the influential work by Owens et al. [Owens et al., 2009; Sewell et al., 2010a]. This formalisation is operational, meaning described by way of individual steps of the system, and as such is often easier to follow than other types of semantics [Jones and Astarte, 2018]. We also choose it as it is based around a microarchitectural feature, and as such provides a contrast to axiomatic descriptions, which purposefully abstract away from such considerations. In this formalisation, weak memory effects (specifically, loads appearing to take effect before earlier stores) emerge by the use of a write buffer in each core (acting similarly to a pipeline (Section 2.1) or as part of the cache system [McKenney, 2010]). We first give an operational semantics for a standard sequential language, as if instructions were executed in the order specified by the program text; after extending that semantics with extra rules covering write buffers the weaker behaviours of x86 (Total Store Order) become apparent. More complex formalisations based on microarchitectural mechanisms are outlined in Section 4.4; see also Section 6.1.
@@ 4.3. Write Buffers for x86
We now augment the system of Sections 4.1 and 4.2 with an explicit write buffer for each core, following [Owens et al., 2009]. A write buffer, of type , is a list of pairs , where is a shared variable and is a value being written to , representing a pending write of to . The head (left-hand side) of the write buffer is the oldest write, with new writes being appended to the end (right-hand side) of the write buffer. A core now contains a write buffer, in addition to the local state and command; the system-level structure is otherwise unchanged.
@@ 4.4. Operational Descriptions of More Complex Models
Below, we outline how some of these more complex aspects of weak memory may be approached through an explicit model of a hardware pipeline as described in Section 2.1. However, a focus on microarchitectural details may become cumbersome to work with formally. For this reason many researchers focus on formal representations that abstract away from microarchitectural mechanisms, and focus instead on key relationships between operations across multiple processes in the system. This is typically referred to as an axiomatic approach, to which we turn our attention in Section 5. Operationally one of the best-known approaches is the Promising Semantics [Kang et al., 2017; Lee et al., 2020; Pulte et al., 2019], as discussed in Section 6.2, which combines an operational semantics with global axioms.
@@
In this case we assume all instructions, not just writes, are fetched into the pipeline (generalising the buffer), similarly to (x86-write-buf). Provided that there are no writes to (of the form ), reads of (of the form ), or fences in the section of the pipeline preceding the write in question (), then the write can be issued. A similar rule can be constructed for reads. These constraints encode those of Definition 2.1, modified by fences as in Example 2.2, and allow the weaker behaviours shown in Example 2.1. A more complete operational description of pipelines that conform to the Arm memory model, and includes branches and jumps, is given in [Colvin and Su, 2025].
@@
The release/acquire concepts [Gharachorloo et al., 1990], as introduced in (1) and Example 2.2, can also be described operationally. Continuing the generalisation of write buffers to pipelines, an operation with release semantics can be executed only when at the head of the pipeline; i.e., when all other prior operations (w.r.t the program order) have completed. A load with acquire semantics prevents later memory operations from completing earlier; hence we extend the constraints of (Arm-write-mem) to capture this possibility.
Rule (release-write) is similar to (x86-write-mem) in that a release write is issued only when at the head of the pipeline. Rule (RA-write-mem) extends (Arm-write-mem) with an extra condition, which we highlight with a box. This additional constraint means that an earlier acquire load of any variable prevents a later write from proceeding. For the stronger form of release/acquire consistency in [Gharachorloo et al., 1990] an acquire load additionally requires the absence of release stores in the earlier pipeline, which can be described syntactically similarly to the boxed constraint of (RA-write-mem).
@@
Instruction types such as swap, compare-and-swap, fetch-and-increment, etc., combine read and write operations atomically. These can be described using the concept of a global lock which prevents a core from performing a memory operation while another core holds the lock [Sewell et al., 2010b]. For instance, the atomic swap instruction, , updates the shared variable to hold the value in register , and updates to the initial value of ; this instruction can be defined as follows.
@@ 5.1. Fundamental Relations
| po | program order | Def. 5.2 | sb | sequenced-before | [Batty et al., 2011] |
@@
| co | coherence order | Def. 5.5 | mo | modification order | [Batty et al., 2011] |
| ws | write serialization | [Alglave, 2012] | |||
@@
On the name “reads-from”, here we follow the recent work by Arm [Alglave et al., 2021], where a typical pair means “ reads from ”, with occurring before . Some may find it more intuitive to think of the relation rf as writes-to. Indeed, means that writes to the value read by .
The relations po, co and rf are the “base relations”. The set of memory events and these base relations constitute an “execution graph”, also known as a “candidate execution” [Alglave et al., 2021].
@@
The extended coherence order – defined to be the transitive closure of com [Lahav et al., 2017].
@@ 5.2. Axiomatic Definition of Sequential Consistency
A memory model is said to satisfy the coherence property if all possible executions of any execution graph is coherent. The coherence property is also known as the uniproc property [Alglave, 2012], SC per location [Alglave et al., 2014b], SC per variable [Alglave and Cousot, 2017] or internal visibility [Alglave et al., 2021].
@@ 5.4. Axiomatic Descriptions of More Complex Models
For illustrative purposes, we have presented how axiomatic formalisms can define the straightforward SC memory model and a simplified version of TSO. In practice, axiomatic formalisms have been used to specify a wide range of behaviours related to more complex memory models and instruction types, e.g., [Alglave et al., 2014b; Pulte et al., 2017; Alglave et al., 2021], as surveyed in Section 6.3. Below we give an indication of how these more complex problems are tackled axiomatically.
@@
Now, a pair belongs to the relation precisely when and are interleaved by a write issued by another thread. In other words, means that and are not atomic. Hence, to ensure that every pair in rmw is atomic, we require that no pair in rmw is in , that is, an execution graph enforces the RMW instructions if . Further details and mechanised tool support can be found in the online repository of the Herd tool [Arm Ltd., 2024; Alglave et al., 2021].
@@ 6.1. Mechanistic Formalisms
Mechanistic formalisms describe the effects of weak memory models with structures that operate in the manner of hardware mechanisms. The best-known of such mechanisms is the local write buffers elucidated in Section 4. Local write buffers describe the behaviour of TSO [Sindhu et al., 1992], on which the x86 memory model is based [Owens et al., 2009; Sarkar et al., 2009; Sewell et al., 2010a]. As local write buffers are relatively straightforward to encode, they often serve as the foundation for verification under TSO. Namely, when each write buffer is modelled as a separate process, one can analyse the entire system under Sequential Consistency using existing tools [Burckhardt and Musuvathi, 2008; Travkin et al., 2013; Travkin and Wehrheim, 2014; Wehrheim and Travkin, 2015; Travkin and Wehrheim, 2016; Linden and Wolper, 2010, 2011, 2013; Štill et al., 2016].
The mechanistic formalism of Boudol and Petri [2009] is also based on write buffers that are not local but potentially hierarchical, making this formalism capable of expressing behaviours weaker than those of TSO. Another recent variant of write buffers is read buffers, proposed by Abdulla et al. [2018] to improve various aspects of verification under TSO. Meanwhile, two other major hardware platforms – Power [Sarkar et al., 2011] and Arm [Flur et al., 2016; Pulte et al., 2017] – also have their own mechanistic formulations, albeit much more complex than x86-TSO and harder to encode for verification tools. A semantics based directly on pipelined and speculative execution, as outlined in Section 2.1, has been developed for Arm, RISC-V and x86 processors [Colvin and Su, 2025].
@@ 6.2. General Operational Formalisms: Timestamped Messages, Promises and Views
Hardware-like mechanisms are insufficient for describing the behaviours of compilers – the source of complexity for the C memory model especially. In order to reason about the C memory model operationally, many formulations make use of the notions of timestamped messages and views. Under weak memory models, a write does not become visible to other threads immediately. Thus, each write can be seen as a message that is broadcast by the issuing process, and it is up to the memory model that dictates when this message becomes visible to which other processes. Each process has a view of a subset of all the messages that have been sent out, and each message has a timestamp attached. By looking at the timestamps, a process updates its view according to the memory model. For instance, instead of just , a write-event is now a pair , where is the time at which the event took place. The timestamp can be some future time, meaning that any subsequent read of with return-value must occur at some later time , i.e., . An early use of timestamps is provided in [Podkopaev et al., 2016]; we look at some other prominent formalisations in this vein below.
@@
In this framework, introduced in Kang et al. [2017], each timestamped message is a promise to fulfil a later write. This semantics was later simplified by Lee et al. [2020], and adapted to Arm and RISC-V by Pulte et al. [2019]. The semantics has been used as the basis for verification of non-trivial systems [Tao et al., 2021].
A characteristic of the Promising Semantics is that it is a multi-execution formalism [Moiseenko et al., 2022], as opposed to a single-execution formalism. In a single-execution formalism (e.g., that of Section 4), one behaviour, e.g., a trace of memory operations, corresponds to one execution of the code. In the Promising Semantics a single behaviour comprises more information that just the sequence memory events, and this extra information also covers multiple possible executions of the code. The multi-execution characteristic can also be applied to partial-order-based representations (see Section 6.4).
@@
Formalisms with timestamped messages and similar notions have since been developed on several fronts, for example, Doherty et al. [2019], with subsequent works leading to program logics [Dalvandi et al., 2020; Wright et al., 2021; Doherty et al., 2022; Dalvandi and Dongol, 2022; Dalvandi et al., 2022; Bargmann and Wehrheim, 2023; Wehrheim et al., 2023]. Lahav and Boker [2022] introduced the notion of potentials. In contrast to timestamped messages which reason about stores that have occurred in the past, potentials reason about possible stores in the future.
Timestamped messages, views, and potentials enable general frameworks for handling a wide range of weak memory effects, including the lack of multicopy atomicity (Section 2.4) and out-of-thin air behaviour (Section 3.3). This flexibility is not without disadvantages, however: these formalisms are perhaps the most complex of those surveyed, with the details of even small examples such as the store buffering litmus test being difficult to reason about [Moiseenko et al., 2022].